OpenStudy (anonymous):

Write the expression as either the sine, cosine, or tangent of a single angle. cosine of pi divided by three times cosine of pi divided by five plus sine of pi divided by three times sine of pi divided by five.

OpenStudy (mertsj):

\[\frac{\cos \pi}{3\cos \pi}\div \frac{5+\sin \pi}{3\sin \frac{\pi}{5}}\]

OpenStudy (mertsj):

Is that the problem?